OpenStudy (anonymous):

can some one help me with this problem 2/5 of what number is 5/6?

Parth (parthkohli):

of -> multiply.

hartnn (hartnn):

let 2/5 of 'x' be 5/6 so u can write (2x/5) = 5/6 can u solve further ?

OpenStudy (anonymous):

noooooooooooooo [and i dont get it]

OpenStudy (anonymous):

\[\frac{ 2x }{ 5 }=\frac{ 5 }{ 6 }\]

OpenStudy (anonymous):

Nw Cross Multiply... 12x = 25 x = 25/12

OpenStudy (anonymous):

okay